OpenStudy (anonymous):

Horatio has a pack of candies containing 4 strawberry, 2 lemon, 1 lime, and 3 orange flavored chews, in random order. If he eats each candy immediately after taking it out of the pack, which multiplication expression can be used to determine the probability that the first two candies he eats are lemon?

Parth (parthkohli):

P(Lemon at the first time) = 2/10 = 1/5 P(Lemon at the second time) = 1/9 Multiply both 1/5 * 1/9 1/45